BexleyBrooks:

The larger number is three less than twice the smaller number. If the sum of the two numbers is 12, find both numbers

AlreadyWinning:

i dunno bud

AcornTheNut:

then why reply?

Ray11:

9 and 3 cuz 9 - 3 = 6 6/2 = 3 3 + 9 = 12 Right??

ykbatman:

wait what

surjithayer:

let the numbers be x and y y>x x+y=12 y=12-x y=2x-3 so 2x-3=12-x 2x+x=12+3 3x=15 \[x=\frac{ 15 }{ 3 }\] x=5 \[y=2\times5-3=7\]